What is the area of a rectangle with width 10x, the length is 3x - 74 is 2500. Please show me how you came to this answer. Thanks, elaine. I am confused

Answer by checkley79(3341)

A=L*W
2500=10X*(3X-74)
2500=30X^2-740X
30X^2-740X-2500=0
10(3X^2-74X-250)=0
x = (-b +- sqrt( b^2-4*a*c ))/(2*a)
X=(+74+_SQRT(-74^2-4*3*-250))/2*3
X=(+74+-SQRT(5476-4*3*-250))/6
X=(74+-SQRT(5476+3000))/6
X=(74+-SQRT(8476)/6
X=(74+_92.065)/6
X=74+92.065)/6
X=166.065/6
X=27.677 ANS.
X=(74-SQRT(8476)/6
X=(74-92.065)/6
X=(-17.065)/6
X=-2.844 ANS.